at 27 degree c, 500ml of helium  diffuse inj 30 min. what is the time taken for 100ml of so2 to diffuse under same experimental conditions

Grade:9

1 Answers

Arun
4 years ago
From Graham's law of diffusion , the rate of diffusion of a gas is inversely proportional to the square root of its molar mass.
⇒ r ∞ 1/ √M
⇒ r = k/√M        where , k is a constant
Also r = V/t  i/e rate of diffusion of volume of gas
Now ,
Molar mass of He = M₁ = 4 gm
Volume = V₁ = 500 ml 
t₁ = 30 mins
Molar mass of SO₂ = M₂ = 64 gm
Volume = V₂ = 100 ml 
t₂ = ? 
rate of diffusion of Helium = r₁  = V₁/t₁  = 500/30 ml/mins
r₂ = 100/t₂ 
r₁ = k/√M₁        and      r₂ = k/√M₂
r₁ / r₂ = k/√M₁ / k /√M₂ = 500/30/100/t₂ 
t₂ = 24 mins 
 
Regards
Arun (askIITians forum expert)